OpenStudy (anonymous):

Solve this system of equations using any method. 4x - y = 5 x + y = 10

OpenStudy (anonymous):

y = 4x - 5 y = mx + c where "m" is the slope

OpenStudy (anonymous):

slope = 4

OpenStudy (mathstudent55):

Use elimination. Add the equations: 5x = 15 x = 3 Now substitute the value of x into original second eq: 3 + y = 10 y = 7

OpenStudy (anonymous):

thank you that helps

OpenStudy (anonymous):

x=3, y=7